Given that g(x)= x^3, find the following: g(2+h)
Answer by Alan3354(69443)

Given that g(x)= x^3, find the following: g(2+h)
---------------
Sub (2+h) for x
g(2+h) = (2+h)^3 = 8 + 12h + 6h^2 + h^3